b. Find the area of a trapezoid with bases of 9 centimeters and 7 centimeters, and a height of 3 centimeters.
Answer by venugopalramana(3286)

FORMULA FOR AREA OF A TRAPEZIUM IS
A=(1/2)*{SUM OF PARALLEL SIDES}*HEIGHT..........HERE WE HAVE
A=(1/2)(9+7)*3=24 SQ.CM.